OpenStudy (anonymous):

An engineer on the ground is looking at the top of a building the angle of elevation to the top of the building is 22 degrees the engineer knows the building is 450 ft tall. What is the distance from the engineer to the base of the building to the nearest whole foot?

OpenStudy (anonymous):

okay, so first draw a picture. You should have a right triangle, with one side being 450 feet, the other side "x" (you want to find that one) and the angle opposite the 450 foot side should be 22 degrees. Then you need to use a little trigonometry. Sin(θ)= (opposite side)/ hypotenuse, cos(θ)= (adjacent side)/ hypotenuse, and tan(θ)= (opposite side)/ (adjacent side). In this case, you will want to use tan. Simply plug in the values! θ= 22˚, the opposite side= 450 feet, and the adjacent side= "x" So solve tan(22)= (450)/x in other words x= (450)/ tan(22) use a calculator, and there you have it

OpenStudy (anonymous):

would it be 1,818 , 1,201, 1,114, 990

OpenStudy (anonymous):

? I got 1113.789 (remember to use degrees, not radaians when you take the tangent)

OpenStudy (anonymous):

the answer is? idk

OpenStudy (anonymous):

When you round to the whole foot, the answer is 1114 feet. Don't forget to round [x